OpenStudy (anonymous):

could someone help me with geometry? im super behind and dont understand half the stuff they are talking about.

OpenStudy (anonymous):

@chosenmatt could you help me?

OpenStudy (nikato):

What do you want help on???

OpenStudy (mstoldegon):

"Identify a pair of alternate exterior angles" Exterior angles = 1, 2, & 6 - These are "outside" the two lines which may be are are implied to be parallel. 3 & 4 are interior (to the parallel lines) Now look for Alternate angles, that is angles that will be on Alternate sides of the intersecting line. Alternate angles = 2 & 6. Angle 1 is on the same side of the intersecting line as 6, and 1 and 2 are not Exterior and Alternate to each other.

OpenStudy (anonymous):

ok. so interior angles are inside the parallel lines and exterior angles are outside. and alternate would be 2 and 6. how am i doing so far @mstoldegon?

OpenStudy (mstoldegon):

So for, so good.

OpenStudy (anonymous):

ok. well now i have another one.

OpenStudy (mstoldegon):

2 is an interior angle but 6 is not. 2 & 6 are same side but 6 is not an interior angle. Corresponding angles means that the angles have the same relation to the same line or parallel lines. Do you think they are corresponding or do you think None of these is correct?

OpenStudy (anonymous):

corresponding?

OpenStudy (mstoldegon):

Maybe to better clarify corresponding angles, they have the same angle and direction meaning if you copied one of the angle and, without rotating it at all, placed it on the other, it would be the same. Does that make more sense?

OpenStudy (anonymous):

yes. that makes much more sense.

OpenStudy (anonymous):

so 2 and 6 would be corresponding?

OpenStudy (mstoldegon):

Yes. If you were told to find the corresponding angle to 6, the only one that has the same Angle (in degrees) AND has the same orientation, would be 2. Note: Corresponding angles Do Not have to have any lines in common. Using the example of runways, if there were two sets of parallel runways that crossed like the one runway crosses the two parallel runways, then there would be 4 corresponding angles on each of the four corners of each intersection. Also, in addition to 2 & 6 being corresponding angles in your example, 1 & 5; 3 & 7; & 4 & 8 are also corresponding angle pairs. Is this making any more sense?

OpenStudy (anonymous):

yes. it does.

OpenStudy (anonymous):

would you be willing to help me on 16 more problems?

OpenStudy (anonymous):

i know its a lot and its late.

OpenStudy (anonymous):

if you cant i could always get someone elses help tomorrow.

OpenStudy (mstoldegon):

I might be able to if I can keep MY answers short.

OpenStudy (anonymous):

ok. um heres the next one.

OpenStudy (mstoldegon):

Do you recall that when two (straight) lines intersect, that any two adjacent angles will add up to 180 degrees, and that opposite angles are equal? 5 = 2, and 2 and 135 deg are equal (they are corresponding).

OpenStudy (mstoldegon):

Sorry it took so long, helping another also.

OpenStudy (anonymous):

its fine. and yes. so since angle 5 and 8 are adjacent that means they have the same angles?

OpenStudy (mstoldegon):

I don't see angle 8 but I will assume is it the one marked 135 deg. if so: angles 1 and 8 are Opposite, as are 2 and 5 opposite angles. 2 and 8 are Corresponding but Not opposite. Angles 3 and 7 are Adjacent angle to 8 (135 deg), and angles 4 & 6 are Adjacent angles to angles 2 & 5. I want to make sure you understand Adjacent versus Opposite.

OpenStudy (mstoldegon):

,,,2 & 5 are Opposite angles TO EACH OTHER, not 1 or 8 to make that clear.

OpenStudy (anonymous):

oh. ok. but angle 5 is 135 deg. right?

OpenStudy (mstoldegon):

Yes.

OpenStudy (mstoldegon):

66 is the answer. Can you explain why? If not, I'll help.

OpenStudy (anonymous):

the best i can come up with is angle q looks smaller that 76 but bigger than 38.

OpenStudy (mstoldegon):

Ok, I'm back. Sorry. Working a Electric Field problem. The angle just below R is an angle Corresponding to the 38 deg angle, right? If so, that angle is also 38 deg. And if Opposite angles are equal, then the angle between Q and 76 deg is also 38 deg. All angles on one side of a straight line will total up to 180 deg, no matter how many angles there are. In this case, there are three (3). So, 76 + 38 + Q = 180. Or 180-(76+38)=Q.

OpenStudy (anonymous):

oh. ok. that makes sense. i didnt look at it that way.

OpenStudy (mstoldegon):

To start with, the values don't agree with the angles shown (they are backwards at least). It says 'not to scale' but WOW! Anyway, lets think about previous examples. When a line intersected two parallel lines, the angles on each side of the intersection (not the ones that were opposite) would add up to 180 deg. Do you understand that? If we imagine that the line with the two angle values was extended beyond one or both lines, we could use the Corresponding 'equivalency' to know that teh angle of value 5x would also be below the line g and the 5x plus the 9x+26 would add up to 180. So, 5x+9x+26=180. Would do you get for x?

OpenStudy (anonymous):

idk. i tried reversing that but got 70. i have a strong feeling i did it wrong.

OpenStudy (mstoldegon):

OK. One hint about a possible answer. Notice that 11 and -11 are possible. Instructors sometime put the correct value down but either as both a plus AND minus or maybe as 11 and 1.1 . They expect that you might get 11 or 1.1 but working the equation backwards you might reverse the signs. Just a little hint. In this case, 5x+9x+26=180; 14x+26=180; 14x=180-26; 14x=154; x=154/14 and finally, x=11 (you won't have any minus angles here). Make sense?

OpenStudy (anonymous):

yea. i guess it does a little. i actually have to go to bed now so this is the last one. but thank you so so much for staying up this late to help me out. i know i must not have been the easiest person to help out. if i could give you a million medals i would but i can only give one...so here ya go. ;)

OpenStudy (mstoldegon):

No problem. Glad to help.